# Symplectic torus actions on complete intersections

### v1

@@ 2.1 Hamiltonian circle actions
This formula follows from the formula of [12, Section 5.8], expressing the signature as a sum over the fixed point components.
We will need the following result which was proved in [20].
@@
[20, Proposition 3.1, Lemma 3.2] Suppose that $M$ is a closed symplectic manifold of dimension $8k$. Suppose that $M$ has a Hamiltonian $S^{1}$-action such that each component of the fixed point set is an isolated fixed point or has dimension $2\mod 4$. Suppose also that $b_{i}(M)=b_{i}(\mathbb{CP}^{4k})$ for $i\neq 4k$. Then the intersection form on $H^{4k}(M,\mathbb{Z})$ is positive definite.
2.2 Applications of Duistermaat-Heckman
We begin by recalling the Dusitermaat-Heckman theorem.
@@ 2.2 Applications of Duistermaat-Heckman
Suppose that $(M,\omega)$ is a closed symplectic manifold with a Hamiltonian $S^{1}$-action, with Hamiltonian $H$. Let $c$ be a regular value of $H$. Then
$\frac{d}{dt}[\omega_{t}]=e(H^{-1}(M)).$
Here $H^{-1}(M)$ is considered as an (orbifold) principal $S^{1}$-bundle over the reduced symplectic orbifold $M_{t}$. Finally, $e$ denotes the Euler class of this bundle in $H^{2}(M_{t},\mathbb{R})$.
Here we denote by $M_{\min}$ and $M_{\max}$ the fixed point components on which the Hamiltonian obtains its minimum and maximum respectively. Recall that they are both connected, symplectic submanifolds.
We call a fixed point component $Y$ internal if $H(M_{\min})<H(Y)<H(M_{\max})$.
@@
Suppose that $M$ is a closed symplectic manifold with a Hamiltonian circle action. Suppose that $M_{\min}$ consists of an isolated fixed. Suppose that $Y$ is an internal fixed point component such that $H(Y^{\prime})\geq H(Y)$ for any other internal fixed component $Y^{\prime}$. Then, $\lambda_{Y}=1$.
@@
This is an immediate consequence of [29, Lemma 3.1]. ∎
We will need the following result which follows from results of Ono.
Lemma 2.7.
[14, Lemma 5.2] Suppose that $(M,\omega)$ is a closed symplectic manifold with a Hamiltonian $S^{1}$-action and $b_{2}(M)=1$. Then after possibly rescaling the symplectic form by a positive constant we may assume that $c_{1}(M)=[\omega]$.
Proof.
Since $b_{2}(M)=1$, $c_{1}(M)=k[\omega]$ for some $k\in\mathbb{R}$. The result then follows from [25, Theorem 1,Theorem 2].
For Hamiltonian $S^{1}$-actions on Fano varieties, the following normalization for the Hamiltonian goes back to Futaki [11, Theorem 1]. In the case of monotone symplectic manifolds proofs may be found for the semi-free case in [6] and in the general case [14]. In [21, Proposition 1.9] the formula was proven for closed symplectic manifolds with Hamiltonian $S^{1}$-actions, such that $c_{1}$ and $[\omega]$ have the same evaluation on classes in the image of the Hurewicz map $\pi_{2}(M)\rightarrow H_{2}(M)$, this will not be needed in this paper.
Theorem 2.8.
Suppose that $(M,\omega)$ is a closed symplectic manifold with $c_{1}(M)=[\omega]$. Then, after possibly adding a constant to the Hamiltonian, for each fixed component $F$,
$H(F)=-\sum w_{i},$
where $w_{i}$ are the weights along $F$.
There has been a lot of recent progress towards classifying monotone symplectic manifolds with Hamiltonian symmetries, see [4, 2, 14, 21, 28] and the reference contained therein.
The following theorem is due to Ohta and Ona, and builds on fundamental work of Gromov, McDuff and Taubes.
Theorem 2.9.
[26] Let $(N,\omega)$ be a closed symplectic $4$-manifold such that $[\omega]=kc_{1}(N)$ for some $k>0$. Then, $N$ is diffeomorphic to a del Pezzo surface.
The theorem has the following immediate corollary.
Corollary 2.10.
Let $(N,\omega)$ be a closed symplectic $4$-manifold such that $[\omega]=kc_{1}(N)$ for some $k>0$. Then, $b_{2}^{+}(N)=1$.
Proof.
By Theorem 2.9, $N$ is diffeomorphic to a smooth del Pezzo surface. By the classification of del Pezzo surfaces, $N$ is either diffeomorphic to $\mathbb{CP}^{1}\times\mathbb{CP}^{1}$ or the blow up of $\mathbb{CP}^{2}$ in up to $9$ points. One may verify by direct calculation that each of these manifolds satisfies $b_{2}^{+}=1$. ∎
2.3 Complete Intersections
A smooth complete intersection $M=X^{n}\subset\mathbb{CP}^{n+k}$ is a subvariety of projective space cut out by a collection of smooth hypersurfaces in general position. By applying the Lefschetz hyperplane consecutively on the image of the variety via appropriate Veronese embeddings, it follows that complete intersections satisfy $b_{i}(X)=\mathbb{CP}^{n}$ for $i\neq n$. The topological consequences of this fact have a different flavour when $n$ is odd or even. In both cases, there is an important characterization of when a complete intersection is“close to projective space”, which is due to Ewing and Moolgavkar when $n$ is odd and Libgober and Wood when $n$ is even.
In the case when $n$ is odd, the middle degree homology $H^{n}(M,\mathbb{R})$ has the structure of a symplectic vector space given by the intersection product and does not contain cup products of cohomology classes of other degrees. Thus the “odd part” of the cohomology ring of $M$ is essentially determined by $b_{n}(M)\in 2\mathbb{Z}$. A complete characterization of when this part of the cohomology ring vanishes, as it does for projective space, was given by Ewing and Moolgavkar [9]. Let $Q=X_{n}(2)$ denote the quadric $n$-fold.
Theorem 2.11.
[9, Corollary 5] Suppose that $X$ is a complete intersection of odd complex dimension $n$, then if $\chi_{top}(M)=n+1$ then $X\cong\mathbb{CP}^{n}$ or $X\cong Q$.
In the case when $n$ is even, then the intersection form on $H^{n}(M,\mathbb{R})$ is symmetric, and $H^{n}(M,\mathbb{R})$ plays a more interesting role in determining the cohomology ring of $M$. The most fundamental invariant we can extract from the intersection form is the signature $\sigma(M)$. A property satisfied by projective space is that the signature is a equal to the alternating sum of even Betti numbers, as was shown by Libgober and Wood, this almost gives a characterization of projective space amongst complete intersections.
Theorem 2.12.
[19] Let $X$ be a complete intersection with $n$ even. Then
$\sigma(X)=\sum_{k\in\mathbb{Z}}(b_{4k}(X)-b_{4k+2}(X))$
if and only if $X$ is equal to $X_{n}(1),X_{n}(2),X_{2}(3)$ or $X_{n}(2,2)$.
Combining the condition $\sigma(X)=\sum_{k\in\mathbb{Z}}(b_{4k}(X)-b_{4k+2}(X))$ with the Lefschetz hyperplane theorem shows that this is equivalent to the intersection form being positve definite when $4|n$ and $b_{n}^{+}(X)=1$ otherwise.
A beautiful fact, proved by Deligne [7, Page 339] is that when $4|n$ the intersection form of $X_{n}(2,2)$ is the Korkine-Zolotareff form $\Gamma_{n+4}$. That is the lattice of $\mathbb{R}^{n+4}$ generated by $e_{i}+e_{j}$ and $\frac{1}{2}(e_{1}+\ldots+e_{n+4}).$ In particular for $n=4$ the intersection form is isomorphic to the $E_{8}$ lattice. Deligne gives a Dynkin diagram description of the intersection form for $n=2\mod 4$, and it is clear from this description that $b^{+}_{n}(X_{n}(2,2))=1$.
3 Proof of Theorem 1.1
We now proceed to the proof of our main result Theorem 3.4. Before giving the details of the proof, we describe the general approach. It is necessary to make the following definition. We follow the convention that for manifolds with real dimension not divisible by $4$ the signature is 00, and a point has signature $1$.
Definition 3.1.
Suppose that $X$ is a closed manifold. Define
$I_{JR}(X)=\sigma(X)-\sum_{i\in\mathbb{Z}}(b_{4i}(X)-b_{4i+2}(X)),$
we refer to $I_{JR}(X)=0$ to be the JR equation.
The idea of the proof of the following proposition was motivated by a theorem of Jones and Rawnsley [16, Theorem 1.1], which states that if a closed symplectic $(M,\omega)$ has a Hamiltonian circle action with isolated fixed points then it satisfies the JR equation.
In the following Proposition 3.2 we show that the invariant $I_{JR}$ localizes in a simple way over the fixed point set. A form of Proposition 3.2 with additional conditions has been proven by Faber [10, Corollary 6]. Fabers result states that for a manifold with a Hamiltonian circle action such that $I_{JR}(F)=0$, $\forall F\subset M^{S^{1}}$ and all the odd Betti numbers of each $F$ vanish, ones has $I_{JR}(M)=0$. Proposition 3.2 improves on [10, Corollary 6] in three ways. Firstly, it removes the assumption on the odd Betti numbers. Secondly, it shows that only fixed submanifolds of positive dimension divisible by $4$ are relevant. Lastly, the equation of Proposition 3.2 can be useful for an inequality analysis when both sides are not necessarily zero, this is key for the proof of Theorem 5.1.
Proposition 3.2.
Suppose that $(M,\omega)$ is a closed symplectic manifold admitting a Hamiltonian circle action. Then,
Proof.
@@ 2.3 Complete Intersections
By Theorem 2.2 and Theorem 2.3 the alternating sum of even Betti numbers and the signature may be be expressed as sums over the fixed points set, in a similar way to the proof of [20, Proposition 3.1]. By Theorem 2.2:
@@
For each component $F$, considering how shifting index changes the sign of an alternating sum gives the equation:
@@
Using these formulas to express $I_{JR}(M)$ as a sum over the fixed point set, gives the equation.
@@
It remains to show that when $F$ is an isolated fixed point or that $\dim(F)=2\mod 4$, that $I_{JR}(F)=0$. For $F$ an isolated fixed point, the statement is straightforward to prove, recalling that a point has signature $1$.
Suppose $N$ has dimension $2\mod 4$, it follows that $\sigma(N)=0$. So to prove $I_{JR}(N)=0$ it is sufficient to prove
@@
This follows from Poincarè duality and the assumption that $\dim(N)=2\mod 4$.
The following simple lemma simplifies $I_{JR}$ for $4$-dimensional components.
@@ 3.1 Proving the JR equation for $4$-dimensional fixed components
In the below theorem, we complete the main step to classifying $8$-dimensional complete intersections with a symplectic circle action. The most important result which motivates Theorem 3.4 is a result of Tolman [29], which classified Hamiltonian $S^{1}$-actions on $6$-manifolds with $b_{2}=1$. In [13], building on results of [14], substantial progress towards the problem of classifying $8$-dimensional symplectic manifolds having a Hamiltonian torus action with isolated fixed points and $b_{2}=1$ was completed. A sharp classification result into four explicit cases was obtained when the torus is two dimensional and $b_{4}=2$ or equivalently the action has exactly $6$ fixed points [13, Main Theorem]. All of the examples subcircle actions with various fixed components of positive dimension occurring.
@@
Suppose that $(M,\omega)$ is a closed symplectic $8$-manifold admitting a Hamiltonian circle action, and that $b_{2}(M)=1$. Then, the intersection form of $M$ is positive definite.
@@
Firstly we note that the conclusion of the theorem is equivalent to showing that $I_{JR}(M)=0$. By Theorem 3.2 and Lemma 3.3 it is sufficient to prove for each $4$-dimensional fixed point component $N$, $b_{2}^{+}(N)=1$.
Let $N$ be a $4$-dimensional fixed component, since the downward flowing normal bundle of $N$ has rank at most 2 (over $\mathbb{C}$), it follows that $0\leq\lambda_{N}\leq 2$. We claim that if $\lambda_{N}$ is 00 or $2$, then by Theorem 2.2 $b_{2}(N)=1$. Since $N$ is a symplectic submanifold $b_{2}(N)>0$. In the localisation sum for the Betti numbers Theorem 2.2, $b_{2}(N)$ either contributes to $b_{2}(M)$ or $b_{6}(M)$ (when $\lambda_{N}=0$ or $2$ respectively) via the localisation formula, but by the assumption of Theorem 3.4 and Poincaré duality $b_{2}(M)=b_{6}(M)=1$. It follows that $b_{2}(N)=1$ in each case. Finally, since $b_{2}(N)=1$ and $N$ is a symplectic submanifold $\int_{N}[\omega_{N}]^{2}>0$ implying that $b_{2}^{+}(N)=1$, proving Theorem 3.4 in this case.
@@
By Theorem 2.2 every other fixed component is an isolated fixed point $p$ with $\lambda_{p}=2$. We claim that such fixed points dont occur. If $H(p)\leq H(N)$ then this follows from Lemma 2.6. If $H(p)\geq H(N)$ then this follows from Lemma 2.6, the case $H(p)\leq H(N)$ is proved similarly.
Hence, there are three fixed components $M_{\min}$, $M_{\max}$ and $N$. By Lemma 2.7 we may assume that $c_{1}(M)=[\omega]$. Since $\dim(N)=4$ and $\lambda_{N}=1$ the normal bundle of $N$ splits as a sum of two line bundles, one upward flowing with respect to the Hamiltonian, and one downward flowing. We denote the splitting by $L_{1}\oplus L_{2}$. Let $w_{1}$ and $w_{2}$ denote the weights of the circle action along $L_{1}$ and $L_{2}$ respectively.
If either of $w_{1}$ or $w_{2}$ has modulus greater than $1$, there is an isotropy $6$-manifold $Q$ such that $Q^{S^{1}}$ has two fixed components. Suppose without loss of generality that $w_{1}<-1$, so that $Q_{\max}=N$, $Q_{\min}=M_{\min}$. Then applying Theorem 2.2 to the Hamiltonian $S^{1}$-action on $Q$ gives $b_{2}(Q)=b_{0}(N)=1$. Then, by Poincaré duality $b_{4}(Q)=1$, then applying Theorem 2.2 again $b_{4}(Q)=b_{2}(N)=1$. Then since the symplectic form of $N$ squares positively we have that $b_{2}^{+}(N)=1$ as required.
Hence, we may assume that $w_{1}=-1$ and $w_{2}=1$. In particular, $H(N)=0$ by Theorem 2.8. To ease notation we set $H(M_{\max})=H_{max}$ and $H_{\min}=H(M_{\min}).$ Then since $M_{\min}$ and $M_{\max}$ are isolated points, all of the four weights along them are non-zero and of the same sign. This implies by Theorem 2.8 that $H(N)-H_{\min}=-H_{\min}\leq-4$ and $H_{\max}-H(N)=H_{\max}\geq 4$.
@@
$[\omega]|_{N}=c_{1}(M)|_{N}=c_{1}(N)+c_{1}(L_{1})+c_{1}(L_{2}).$ (1)
Let $\bar{N}$ be the union of gradient sphere with maximum in $N$. We consider the cycle $N_{c}\subset M_{c}$ which is obtained as the projection of $\bar{N}$ by the quotient map $N_{c}$. Considering levels of the Hamiltonian slightly below $0=H(N)$, and applying the equivariant symplectic neighborhood theorem gives that:
$e(H^{-1}(c))|_{N_{c}}=c_{1}(N_{1})|_{N_{c}}\in H^{2}(N_{c},\mathbb{Z}).$
Note that $N_{H_{\min}}$ is a point since $M_{min}$ is isolated, by the equivariant Darboux threorem, the limit of the class of the reduced symplectic form is zero as the Hamiltonian approaches $H_{\min}$. Applying Duistermaat-Heckman Theorem 2.5 and the above observations yields the equation:
$[\omega]|_{N}=(H(N)-H_{\min})c_{1}(L_{1})=|H_{\min}|c_{1}(L_{1})$
Applying similar consinderations to the upward flowing manifold of $N$ gives the following equation:
$[\omega]|_{N}=(H_{\max}-H(N))c_{1}(L_{2})=|H_{\max}|c_{1}(L_{2})$

### v2

@@ 2.3 Applications of Duistermaat-Heckman
The first claim follows from applying Lemma 2.7. The second statement follows from applying Lemma 2.7 to the Hamiltonian $S^{1}$-action with generating vector field X-X and Hamiltonian H-H.∎
A theorem of Ono states that is closed symplectic manifold $(M,\omega)$ satisfies $c_{1}=k[\omega]$ $k\in\mathbb{R}$ then if $k\leq 0$ $(M,\omega)$ cannot have a non-trivial Hamiltonian $S^{1}$-action [35, Proposition 1.1]. It was also proved that if $k<0$ that $(M,\omega)$ cannot have a symplectic $S^{1}$-action [35, Theorem 1].
Lemma 2.9.
Suppose that $(M,\omega)$ is a closed symplectic manifold with a Hamiltonian $S^{1}$-action and $b_{2}(M)=1$. Then after possibly rescaling the symplectic form by a positive constant we may assume that $c_{1}(M)=[\omega]$.
Proof.
Since $b_{2}(M)=1$, $c_{1}(M)=k[\omega]$ for some $k\in\mathbb{R}$. Therefore the result follows immediately from [35, Proposition 1.1].∎
For Hamiltonian $S^{1}$-actions on Fano varieties, the following normalization for the Hamiltonian goes back to Futaki [13, Theorem 1]. In the case of monotone symplectic manifolds proofs may be found for the semi-free case in [7] and for the general case in [17].  [^1] In [30, Proposition 1.9] the formula was proven for closed symplectic manifolds with Hamiltonian $S^{1}$-actions, such that $c_{1}$ and $\omega]$ have the same evaluation on classes in the image of the Hurewicz map $\pi_{2}(M)\rightarrow H_{2}(M)$, this will not be needed in this paper.
@@
Suppose that $(M,\omega)$ is a closed symplectic manifold with $c_{1}(M)=[\omega]$ with a Hamiltonian $S^{1}$-action with Hamiltonian $H:M\rightarrow\mathbb{R}$. Then, after possibly adding a constant to the Hamiltonian, for each fixed component $F$,
$H(F)=-\sum_{i=1}^{n}w_{i},$
where $w_{i}$ are the weights along $F$.
The following theorem is due to Ohta and Ono, and builds on fundamental work of Gromov, McDuff and Taubes.
@@
[36] Let $(N,\omega)$ be a closed symplectic $4$-manifold such that $\omega]=kc_{1}(N)$ for some $k>0$. Then, $N$ is diffeomorphic to a del Pezzo surface.
The theorem has the following immediate consequence.
Corollary 2.12.
Let $(N,\omega)$ be a closed symplectic $4$-manifold such that $\omega]=kc_{1}(N)$ for some $k>0$. Then, $b_{2}^{+}(N)=1$.
Proof.
By Theorem 2.11, $N$ is diffeomorphic to a smooth del Pezzo surface. By the classification of del Pezzo surfaces, $N$ is either diffeomorphic to $\mathbb{CP}^{1}\times\mathbb{CP}^{1}$ or $\mathbb{CP}^{2}\#k\overline{\mathbb{CP}^{2}}$, where $0\leq k\leq 8$. One may verify by direct calculation that each of these manifolds satisfies $b_{2}^{+}=1$.∎
2.4 Complete Intersections
We start this subsection by recalling a basic definition of algebraic geometry. Let $X\subset\mathbb{CP}^{n+k}$ be a smooth subvariety of complex dimension $n$.
Definition 2.13.
$X$ is called a complete intersection if it is a transversal intersection of $k$ smooth hypersurfaces.
Since we restrict to the smooth case transversality can be taken simply in the sense of complex submanifolds, but also has algebraic reformalizations that are useful; such as in terms of gradients or the ideal associated to $X$. Let $\{d_{i}|\;i=1,\ldots,k\}$ be the multiset consisting of the degrees of the $k$ hypersurfaces. Let $X_{m}(d_{1},\ldots,d_{k})$ denote a smooth complete intersection in $\mathbb{CP}^{k+m}$ with multidegree $(d_{1},\ldots,d_{k})$. A well-known consequence of Ehresmann’s Theorem is that the diffeomorphism type of $X_{m}(d_{1},\ldots,d_{k})$ depends only on the dimension and the multidegree. Hence as a smooth manifold the notation $X_{m}(d_{1},\ldots,d_{k})$ is unambiguous. Note however, that complete intersections with different multidegree can sometimes be diffeomorphic, a simple example being furnished by $X_{1}(3)\cong X_{1}(2,2)\cong S^{1}\times S^{1}$. Another basic fact following from the adjunction formula is that a complete intersection is Fano resp. generalized Calabi-Yau resp. general type, when $\sum_{i}d_{i}$ is less than resp. equal to resp. greater than $n+1$.
By applying the Lefschetz hyperplane theorem consecutively on the image of the variety via appropriate Veronese embeddings, the following holds:
@@ 2.4 Complete Intersections
Let $X$ be a smooth complete intersection of complex dimension $n$, then $b_{i}(X)=b_{i}(\mathbb{CP}^{n})\text{ for }i\neq n.$
The topological consequences of this fact have a different flavor when $n$ is odd or even. In both cases, there is an important characterization of when a complete intersection is “close to projective space”, which is due to Ewing and Moolgavkar when $n$ is odd [11, Corollary 5] and Libgober and Wood when $n$ is even [28, Corollary 6.1].
In the case when $n$ is odd, the middle degree cohomology $H^{n}(M,\mathbb{R})$ has the structure of a symplectic vector space given by the intersection product and does not contain cup products of cohomology classes of other degrees. Thus the “odd part” of the cohomology ring of $M$ is essentially determined by $b_{n}(M)\in 2\mathbb{Z}$. A complete characterization of when this part of the cohomology ring vanishes, as it does for projective space, was given by Ewing and Moolgavkar [11].
Theorem 2.15.
[11, Corollary 5] Let $X$ be a complete intersection with odd complex dimension $n$, then $\chi(M)=n+1$ if and only if $X$ is diffeomorphic to $X_{n}(1)$ or $X_{n}(2)$.
In the case when $n$ is even, then the intersection form on $H^{n}(M,\mathbb{R})$ is symmetric, and $H^{n}(M,\mathbb{R})$ plays a more interesting role in determining the cohomology ring of $M$. The most fundamental invariant we can extract from the intersection form is the signature $\sigma(M)$. A property satisfied by projective space is that the signature is equal to the alternating sum of even Betti numbers, as was shown by Libgober and Wood, this almost gives a characterization of projective space among complete intersections.
Theorem 2.16.
[28, Corollary 6.1] Let $X$ be a complete intersection with even complex dimension $n$. Then
$\sigma(X)=\sum_{k\in\mathbb{Z}}(b_{4k}(X)-b_{4k+2}(X))$
if and only if $X$ is diffeomorphic to $X_{n}(1),X_{n}(2),X_{2}(3)$ or $X_{n}(2,2)$. That is:
  • a.)
    If $4|n$, the intersection form is positive definite if and only $X$ is diffeomorphic to $X_{n}(1),X_{n}(2)$ or $X_{n}(2,2)$
  • b.)
    If $n=2\mod 4$, $b_{n}^{+}(X)$=1 if and only $X$ is diffeomorphic to $X_{n}(1),X_{n}(2),X_{n}(2,2)$ or $X_{2}(3)$
The proof of the equivalence of a.)-b.) with the first statement, follows from writing $\sigma(X)=b_{n}^{+}(X)-b_{n}^{-}(X)$, $b_{n}(X)=b_{n}^{+}(X)+b_{n}^{-}(X)$ and then applying Lemma 2.14.
Remark 2.17.
A beautiful fact, proved by Deligne [8, Page 339] is that when $4|n$ the intersection form of $X_{n}(2,2)$ is the Korkine-Zolotareff form $\Gamma_{n+4}$. That is the lattice of $\mathbb{R}^{n+4}$ generated by $e_{i}+e_{j}$ and $\frac{1}{2}(e_{1}+\ldots+e_{n+4}).$ In particular for $n=4$ the intersection form is isomorphic to the $E_{8}$ lattice. Deligne gives a Dynkin diagram description of the intersection form for $n=2\mod 4$, and it is clear from this description that $b^{+}_{n}(X_{n}(2,2))=1$.
Finally, we state one of the major motivating results of this article due to Benoist. Let $X$ be a smooth projective variety, by a well-known relative version of Chow’s theorem the group of bihomomorphisms of $X$ coincides with the group of isomorphisms of $X$, we denote it by $Aut(X)$.
Theorem 2.18.
[1, Theorem 3.1] Let $X$ be a complex smooth complete intersection of dimension at least two. Then $|Aut(X)|=\infty$ if and only $X$ is isomorphic to $X_{n}(1)$ or $X_{n}(2)$.
There are two remarks to make here. In dimension $1$ a smooth elliptic curve (realized either a cubic plane curve or an intersection of two quadrics in $\mathbb{P}^{3}$) has infinite automorphism group hence the omission of dimension $1$ from Benoist’s theorem. Secondly, as was noted in the introduction, the smooth manifolds underlying complete intersections $X_{2}(2,2)$ and $X_{2}(3)$ do admit algebraic structures with infinite automorphism groups, however these algebraic structure cannot be isomorphic to a complete intersection. To see this, note by the adjuction formula such complete intersections are Del Pezzo, of degree $3$ and $4$ respectively, the finiteness of the automorphism groups of such surfaces was known classically.
3 Proof of Theorem 1.2
We now proceed to the proof of our main result Theorem 3.6. Before giving the details of the proof, we describe the general approach. Firstly we prove Proposition 3.2 which is a localization formula for a certain topological invariant $I_{JR}$ which is defined below. The idea of Proposition 3.2 proposition was motivated by a theorem of Jones and Rawnsley [23, Theorem 1.1], [^2] We note that [23, Theorem 1] was known much earlier for smooth complex projective varieties having an algebraic torus action with only isolated fixed points. It may be proved by combining the Hodge index theorem, and the fact due to Białynicki-Birula that the Hodge numbers such a variety are pure i.e. $h^{ij}=0$ for $i\neq j$ [2]. In this sense [23, Theorem 1] may be considered as a symplectic version of the Hodge index theorem. which states that if a closed symplectic $(M,\omega)$ has a Hamiltonian $S^{1}$-action with isolated fixed points then it satisfies the JR equation. It is necessary to make the following definition.
Definition 3.1.
Suppose that $X$ is a closed manifold. The JR equation is
$I_{JR}(X):=\sigma(X)-\sum_{i\in\mathbb{Z}}(b_{4i}(X)-b_{4i+2}(X))=0.$
In Proposition 3.2 below, it is shown that the invariant $I_{JR}$ localizes in a simple way over the fixed point set.
Proposition 3.2.
Suppose that $(M,\omega)$ is a closed symplectic manifold admitting a Hamiltonian $S^{1}$-action. Then,
$I_{JR}(M)=\sum_{\{F\subset M^{S^{1}}|\dim(F)=4k>0\}}(-1)^{\lambda_{F}}I_{JR}(F).$ (1)
@@
By Theorems 2.2 and 2.3 the alternating sum of even Betti numbers and the signature may be be expressed as sums over the fixed points set, in a similar way to the proof of [29, Proposition 3.1]. By Theorem 2.2:
@@
For each component $F$, by shifting the indices in the above alternating sum, we obtain that:
@@
Hence,
@@
It remains to show that if $F$ is an isolated fixed point or $\dim(F)=2\mod 4$, then $I_{JR}(F)=0$. For $F$ an isolated fixed point, the statement is straightforward to prove, recalling that a point has signature $1$.
Suppose $N$ has dimension $2\mod 4$, by definition $\sigma(N)=0$. Hence, it suffices to prove $I_{JR}(N)=0$ it is sufficient to prove
@@
This follows from Poincarè duality and the assumption that $\dim(N)=2\mod 4$.∎
Remark 3.3.
A form of Proposition 3.2 with additional conditions has been proven by Farber [12, Corollary 6]. Farber’s result states that for a manifold with a Hamiltonian $S^{1}$-action such that $I_{JR}(F)=0$, $\forall F\subset M^{S^{1}}$ and all the odd Betti numbers of each $F$ vanish, it holds that $I_{JR}(M)=0$. Proposition 3.2 improves on [12, Corollary 6] in three ways. Firstly, it removes the assumption on the odd Betti numbers. Secondly, it shows that only fixed submanifolds of positive dimension divisible by $4$ are relevant. Lastly, the equation of Proposition 3.2 can be useful for an inequality analysis when both sides are not necessarily zero, this is key for the proof of Theorem 1.6.
Given the generality of Proposition 3.2, it is natural ask whether $I_{JR}(M)=0$ for all closed symplectic manifolds admitting a Hamiltonian $S^{1}$-action. The following example shows that this is not the case.
Example 3.4.
Consider the standard Hamiltonian $S^{1}$-action on $\mathbb{CP}^{4}$ with the Fubini-Study form, given by
$\lambda\cdot[z_{0}:\ldots:z_{4}]=[\lambda z_{0}:z_{1}:\ldots:z_{4}].$
Then, $(\mathbb{CP}^{4})^{S^{1}}$ has two components $1:0:\ldots:0]$ and the hyperplane $H$ defined by the equation $z_{0}=0$. Let $S\subset H$ be a smooth degree $4$ hypersurface. Recall that $S$ is a K3 surface with signature $16$ and $b_{2}(S)=22$. Let $M$ be the $S^{1}$-equivariant blow-up of $\mathbb{CP}^{4}$ along $S$. Then $M^{S^{1}}$ has three components diffeomorphic to $H$, $S$ and point respectively, moreover the fixed point component diffeomorphic to $S$ satisfies $\lambda_{S}=1$. By Proposition 3.2
$I_{JR}(M)=-I_{JR}(S)=-36.$
The following result simplifies $I_{JR}$ for $4$-dimensional components.
@@ 3.1 Proving the JR equation for $4$-dimensional fixed components
In the theorem below, we complete the main step to classifying $8$-dimensional complete intersections with a symplectic $S^{1}$-action. The most important result which motivates Theorem 3.6 is a result of Tolman [39], which classified Hamiltonian $S^{1}$-actions on $6$-manifolds with $b_{2}=1$. In [15], building on results of [17], substantial progress towards the problem of classifying $8$-dimensional symplectic manifolds having a Hamiltonian torus action with isolated fixed points and $b_{2}=1$ was completed. A sharp classification result into four explicit cases was obtained when the torus is two dimensional and $b_{4}=2$ or, equivalently, when the action has exactly $6$ fixed points [15, Main Theorem]. All of the examples have various subgroup $S^{1}$-actions with various fixed components of positive dimension.
@@
Suppose that $(M,\omega)$ is a closed symplectic $8$-manifold admitting a Hamiltonian $S^{1}$-action, and that $b_{2}(M)=1$. Then the intersection form of $M$ is positive definite.
@@
Firstly we note that the conclusion of the theorem is equivalent to showing that $I_{JR}(M)=0$ (see Theorem 2.16). Moreover, by Theorem 3.2 and Lemma 3.5 if $b_{2}^{+}(N)=1$ for each $4$-dimensional fixed point component $N$, then the result holds.
Let $N$ be a $4$-dimensional fixed component, by a dimension count it follows that $0\leq\lambda_{N}\leq 2$. Suppose that $\lambda_{N}\in\{0,2\}$, then by Theorem 2.2 $b_{2}(N)=1$. Since $N$ is a closed symplectic submanifold $b_{2}(N)>0$. In the localization sum for the Betti numbers Theorem 2.2, $b_{2}(N)$ either contributes to $b_{2}(M)$ or $b_{6}(M)$ (when $\lambda_{N}=0$ or $2$ respectively). By hypothesis and Poincaré duality we have that $b_{2}(M)=b_{6}(M)=1$. It follows that $b_{2}(N)=1$ in each case. Finally, since $b_{2}(N)=1$ and $N$ is a closed symplectic submanifold $\int_{N}[\omega_{N}]^{2}>0$ implying that $b_{2}^{+}(N)=1$.
@@
By Theorem 2.2 every other fixed component is an isolated fixed point $p$ with $\lambda_{p}=2$. By Lemma 2.8 that such fixed points do not occur. More precisely if such fixed points satisfy $H(p)\leq H(N)$ then they are excluded by the first part of Lemma 2.8 and if $H(p)\geq H(N)$ they are excluded by the second part.
Hence, there are three fixed components $M_{\min}$, $M_{\max}$ and $N$. By Lemma 2.9 we may assume that $c_{1}(M)=[\omega]$. Since $\dim(N)=4$ and $\lambda_{N}=1$ the normal bundle of $N$ splits as a sum of two line bundles, one upward flowing with respect to the Hamiltonian, and one downward flowing. We denote the splitting by $L_{1}\oplus L_{2}$. Let $w_{1}$ and $w_{2}$ denote the weights of the $S^{1}$-action along $L_{1}$ and $L_{2}$ respectively.
If either $w_{1}$ or $w_{2}$ has modulus greater than $1$, then there is an isotropy $6$-manifold $Q$ such that $Q^{S^{1}}$ has two fixed components. Suppose without loss of generality that $w_{1}<-1$, so that $Q_{\max}=N$, $Q_{\min}=M_{\min}$. Then applying Theorem 2.2 to the Hamiltonian $S^{1}$-action on $Q$ gives $b_{2}(Q)=b_{0}(N)=1$. Then, by Poincaré duality $b_{4}(Q)=1$, then applying Theorem 2.2 again $b_{4}(Q)=b_{2}(N)=1$. Then since the symplectic form of $N$ squares positively we have that $b_{2}^{+}(N)=1$ as required.
Hence, we may assume that $w_{1}=-1$ and $w_{2}=1$. By Theorem 2.10, $H(N)=0$ To ease notation we set $H(M_{\max})=H_{\max}$ and $H_{\min}=H(M_{\min}).$ Then since $M_{\min}$ (respectively $M_{\max}$) is an isolated fixed point, all of the four weights along it are non-zero and positive (respectively negative). This implies by Theorem 2.10 that $|H_{\min}|=-H_{\min}\geq 4$ and $|H_{\max}|=H_{\max}\geq 4$.
@@
$\omega]|_{N}=c_{1}(M)|_{N}=c_{1}(N)+c_{1}(L_{1})+c_{1}(L_{2}).$ (2)
Next, we use Lemma A.2 to analyse the behavior of the symplectic form around $N$. Let $M_{-}$ denote the symplectic manifold provided by Lemma A.2.a). Then since $H(N)=0$ and $\mu_{-}$ is an equivariant symplectomorphism below level 0 by applying Lemma A.1.
$\omega_{-}]|_{F_{-}}=|H_{\min}|c_{1}(H_{-}^{-1}(0))_{F_{-}}.$
Therefore, by Lemma A.2.c):
$\omega]|_{N}=|H_{\min}|c_{1}(L_{1}).$
Applying the same idea to $M_{+},\mu_{+}$ etc. give by Lemma A.2 gives the equation:
$\omega]|_{N}=|H_{\max}|c_{1}(L_{2}).$